OpenStudy (anonymous):

Find the polynomial of degree 4 whose graph goes through the points (-3, -188), (-2, -35), (0, 7), (1, 16), and (2, 17).

OpenStudy (anonymous):

(-3, -188), (-2, -35), (0,7), (1, 16), and (2, 17) there we go lol

zepdrix (zepdrix):

Hmmmm

zepdrix (zepdrix):

I'm not sure if this is the best approach, but here's one idea... Since it passes through (0,7), it has a y-intercept of 7.\[\Large\bf y=ax^4+bx^3+cx^2+dx+7\]

zepdrix (zepdrix):

We can setup a system of equations using the other points provided.\[\Large (1,16):\]\[\Large\bf 16=a+b+c+d+7\]Then another point gives us:\[\Large\bf (2,17):\]\[\Large \bf 17=a2^4+b2^3+c2^2+d2+7\]

zepdrix (zepdrix):

And so on :o